OpenStudy (keana):

multiply √3*√5 a:5 √3 b:√15 c:3 √5 d:√8

OpenStudy (anonymous):

\[\sqrt{a} \times \sqrt{b} = \sqrt{a \times b}\] Use this formula to get your answer...

OpenStudy (anonymous):

sqrt(3*5)=?

OpenStudy (keana):

i think it's 15

OpenStudy (anonymous):

No...

OpenStudy (anonymous):

well its sqrt(15) is that what u meant?

OpenStudy (keana):

yeah